Four coins are tossed.Let A be the random variable representing the number of heads the occur.Find the value of the random variable A

1
Expert's answer
2022-02-23T14:14:26-0500

This is a binomial distribution with "n=4,p=\frac{1}{2}."

"P(A=0)=C_4^0(\frac{1}{2})^0(\frac{1}{2})^4=\frac{1}{16}."

"P(A=1)=C_4^1(\frac{1}{2})^1(\frac{1}{2})^3=\frac{1}{4}."

"P(A=2)=C_4^2(\frac{1}{2})^2(\frac{1}{2})^2=\frac{3}{8}."

"P(A=3)=C_4^3(\frac{1}{2})^3(\frac{1}{2})^1=\frac{1}{4}."

"P(A=4)=C_4^4(\frac{1}{2})^4(\frac{1}{2})^0=\frac{1}{16}."